# GTAvatar: Bridging Gaussian Splatting and Texture Mapping for Relightable and Editable Gaussian Avatars

### v1

@@ Abstract
Recent advancements in Gaussian Splatting have enabled increasingly accurate reconstruction of photorealistic head avatars, opening the door to numerous applications in visual effects, videoconferencing, and virtual reality. This, however, comes with the lack of intuitive editability offered by traditional triangle mesh–based methods. In contrast, we propose a method that combines the accuracy and fidelity of 2D Gaussian Splatting with the intuitiveness of UV texture mapping. By embedding each canonical Gaussian primitive’s local frame into a patch in the UV space of a template mesh in a computationally efficient manner, we reconstruct continuous editable material head textures from a single monocular video on a conventional UV domain. Furthermore, we leverage an efficient physically based reflectance model to enable relighting and editing of these intrinsic material maps. Through extensive comparisons with state-of-the-art methods, we demonstrate the accuracy of our reconstructions, the quality of our relighting results, and the ability to provide intuitive controls for modifying an avatar’s appearance and geometry via texture mapping without additional optimization.
@@ 1 Introduction
Gaussian Splatting based avatars have revolutionized the capture and rendering of digital humans by delivering an unprecedented level of photorealism with real-time rendering capability, enabling new possibilities for reanimating content from video inputs. However, this realism comes with a trade-off: unlike traditional texture-map-based modeling, Gaussian avatars offer little flexibility for intuitive appearance editing. This gap becomes critical in practice. In film and visual effects, artists routinely adjust the smallest details of a face - smoothing skin to create a flawless appearance, removing distracting high-frequency features, or sculpting age and character through wrinkles, scars and bruises. In gaming and virtual production, creators seek the same level of control to personalize avatars with tattoos, makeup, or stylized patterns using artist-friendly tools that operate directly on material texture maps. Without such editing capabilities, even the most realistic avatar remains a fixed reproduction, rather than a medium for creative expression.
This limitation stems from the fact that each Gaussian splat carries its own color or material properties in isolation, without the shared structure that texture maps naturally provide. As a result, the surface lacks the local coherence needed to treat it as a continuous canvas, making it difficult — if not impossible — to apply edits consistently, whether to a small region of the face or to its overall appearance. Conversely, while mesh-based 3D inverse rendering methods [nvdiffrec, flare, spark] inherit a coherent defacto UV-domain fit for artist-friendly edits, they can suffer from topological rigidity when representing high-frequency details and are hampered in their reconstruction capacity due to the limitations of differentiable rasterization when handling complex or translucent geometries (cf. Table 1). Furthermore, naively embedding Gaussians in UV domain leads to discontinuous texture maps that are hard to edit successfully, as witnessed in our results (e.g. Figure 9) and also in [fate]. The authors of FATE [fate] proposed a second U-Net neural baking stage to alleviate this issue in the context of non-relightable avatars.
@@
Following seminal work [gaussianavatars, surfhead], we adopt FLAME-anchored [FLAME] Gaussian splats. Our key observation is that, unlike vanilla 3DGS [3DGS], which operates directly in screen space, the 2DGS variant [2DGS] defines a local splat coordinate frame and computes ray–splat intersections to query kernels. We leverage this property to map ray–splat intersections into the UV domain via an approximate orthographic projection from the splat tangent space to its corresponding FLAME mesh triangle, effectively mapping each splat plane to a continuous UV patch. We devise an efficient method to compute this mapping using a single matrix multiplication per intersection, which is highly important to maintain the method’s efficiency given the very high number of intersections involved in rendering. PBR material attributes are then sampled smoothly from learnable texture maps (albedo, roughness, and specular reflectance) using bilinear filtering. The splat orientations are combined with a residual normal map to produce the final shading normals. We jointly optimize the FLAME parameters, texture maps, 2D Gaussian positional parameters and environment lighting. This enables differentiable expression of outgoing radiance via the Cook–Torrance BRDF [cook1982reflectance] under the split-sum approximation [karis2013real], using splatted geometry and material G-buffers in a deferred shading framework. Finally, we introduce a novel UV regularization that is crucial for maintaining representation integrity by enforcing alignment of UV coordinates of ray–splat intersections for a given ray. Figure 9 underlines the benefits of our UV mapping and regularization compared to a naive UV embedding strategy based on the projection of splat origins.
@@
  • Expressive Gaussian primitives. Spatially varying attributes and normals, in a Phong shading [phong] fashion, as opposed to single values, yield higher reconstruction quality with a smaller total model size than state-of-the-art relightable avatar competition [HRAvatar] (Figure 6(b)). While similar expressiveness can also be achieved with small per-primitive textures (e.g. [gstex, BBSplat, texturedgaussiansenhanced3d, texturesplat, SuperGaussians, GaussianBillboards, HDGS]), our approach uses a compact global map, offering both efficiency and editability.
@@
    • Intrinsic material supervision, such as our albedo regularization with the FLAME albedo model that guides physical decomposition mid-training and reduces artifacts (Figure 9). We note that we also incorporate a diffusion-based albedo prior in screen space, similar to [HRAvatar], to further disambiguate the decomposition.
@@
We demonstrate the benefits of our method compared to traditional splat-based reconstruction and rendering techniques, and illustrate its versatility through various texture editing examples. Our method produces highly realistic head reconstructions that not only match or surpass the quality of state-of-the-art monocular relightable avatar HRAvatar [HRAvatar] — despite its lack of explicit texture representations — but also unlocks new possibilities for artistic control.
@@ 2 Related work
Avatar Representations. The prior success of NeRFs [nerf, ingp] brought impressive levels of photorealism in neural avatars based on differentiable volume rendering [headnerf, nerface, adnerf]. Rasterization approaches can alleviate the computational rendering burden of implicit neural ray casting. However, mesh-based rasterization [nvdiffrec] can yield subpar photometric reconstruction performance (e.g. FLARE [flare] in Table 1) due to the limitations of meshes in modeling transparent and complex geometry. The recent advent of Gaussian Splatting [3DGS] has enabled state-of-the-art reconstruction quality with real-time rendering [gaussianheadavatar, gavatar, gaussianavatars]. Gaussian Splatting extends the EWA volume resampling framework [ewa, ewa2001] to learnable [adam] inverse rendering, modeling 3D scenes with explicit anisotropic Gaussian kernel primitives that can be sorted and rasterized efficiently in tiles. In this respect, the 2DGS variant [2DGS] leverages planar 2D primitives instead of volumetric ones, and performs precise 2D kernel evaluation in object space as opposed to approximate ones in screen space (3DGS), thus leading to superior geometry and multi-view consistency. The superior normals provided by 2DGS [surfhead, sheap, sparfels, GaussianSurfels] have shown to be beneficial in physics-based inverse rendering [RefGaussian, texturesplat, SVG-IR, GS-2DGS, IRGS] where inferring precise reflection directions [refnerf] is pivotal. Generalizable models for 2DGS [meshsplat, sparsplat] have been recently proposed similarly to 3DGS [pixelSplat, mvsplat] before. 2DGS also defines a local splat coordinate frame that allows us to map ray-splat intersections to our canonical UV space continuously, which is key in our novel avatar representation.
@@
Monocular & Relightable Avatars. Learning avatars from casual consumer grade videos paves the road to democratizing volumetric capture, in contrast to the costly requirements of light stages or multi-view setups [lightStage, relightables, travatar]. Recent advances enable avatar learning from monocular video input (e.g. [pointavatar, flashavatar, monogaussianavatar, splattingavatar, gaussianblendshapes]), thanks to built-in inductive biases and robust monocular facial tracking (e.g. [DECA, EMOCA, SMIRK, MICA]) and reconstruction (e.g. [Pixel3DMM, Sapiens, MonoNPHM, PRN, crossmodal]) methods. Rigging avatars with underlying 3D morphable models (3DMMs) [FLAME, BFM, decoupled, LSFM, CoMA, NonLinear3DMM, facetunegan] enables efficient and robust learning and parametric animation control. They define a canonical space for implicit and explicit representations and factor out expression and pose deformations. Several works leverage data captured under controlled lighting conditions to produce relightable head avatars [rgca, vrmm] with dynamic radiance fields and physical reflectance models. Relightable Gaussian Codec Avatar [Codec, URAvatar] material parameterization has shown good relighting capabilities for full multi-view OLAT captures. Despite the increased ill-posedness of the problem, physics-based inverse rendering and relighting from mere monocular videos or sparse images is possible through further regularizations or appearance priors, such as modeling reflectance via simplified bidirectional reflectance distribution functions (BRDF) [flare, spark], or relying on large data corpora [URAvatar]. Closest to our context, HRAvatar [HRAvatar] employs deferred shading of FLAME-rigged 3D Gaussians augmented with material attributes. In contrast to the latter, we build on 2DGS deferred shading with the aim of extending controllability beyond relighting and reenactment, to additionally include intuitive local editing on conformal material maps, in the interest of more expressive and user-controllable avatar manipulation without sacrificing performance.
@@
UV Mapping & Editing. Recent research has sought to combine explicit surface parameterization with neural representations. Within the NeRF [nerf] framework, works such as NeuTex [xiang2021neutex], Neural Gauge Field [zhan2023general] and Nuvo [srinivasan2024nuvo] learn neural UV mappings that establish bijective correspondences between surfaces and textures, enabling surface-aware rendering across general or category-specific domains. Simultaneously, Gaussian Splatting can support 3D domain editing through semantic grouping and manipulation [gaussian_grouping], dynamic 4D content [yu2024cogs], text-driven scene modifications [chen2024gaussianeditor], or brush painting [painting]. Editing via screen-space gradients [texttoon, mega, PortraitGen], which relies mostly on extensive optimization with pretrained diffusion models [realcompo, zhang2025itercomp] such as Instruct-Pix2Pix [InstructPix2Pix], suffers from high computational costs and lacks precise controllability. UV-domain editing requires continuous texture maps. While some methods structure Gaussians in the UV domain [gghead, GaussianShellMaps, splattingavatar, flashavatar, mega], most still operate with discontinuous sparse maps. Texture-GS [texturegs] learns a neural UV mapping for 3DGS primitives for static objects, whilst we aim to use a predefined UV map for semantically grounded edits of a dynamic avatar. Contemporary work FATE [fate] attempts to alleviate the texture map continuity issue for non-relightable dynamic Gaussian head avatars. However, it requires a separate learning stage involving neural baking via additional networks to obtain smooth editable textures and fails to preserve sharp details when edited, as shown in Figure 11 . Our method avoids these multiple stages, bypassing the need for any neural UV mapping or neural baking. Moreover, to our knowledge, our approach is the first Gaussian head avatar to conjointly enable relighting and UV editing of appearance and normals.
@@ 3.1 FLAME
The FLAME [FLAME] 3D morphable model provides a template head mesh with joints for eyes, jaw and neck, and a statistical model for identity and expressions. Identity and expression deformations are expressed as blendshapes while the joints transform vertices through Linear Blend Skinning [LBS]. For simplicity, we refer to FLAME as template vertex positions $V_{t}\in\mathbb{R}^{3V}$, triangle topology $0..V-1]^{3F}$ and a deformation function $\mathcal{F}$:
@@
where the identity parameters are fixed and $\Psi$ is the concatenation of pose and expression parameters. We refer the reader to [FLAME] for exhaustive details on this 3DMM. Additionally, FLAME provides per-corner UV coordinates for texturing (i.e. UV parameterization for the unwrapped template mesh).
@@ 3.2 Gaussian Splatting
3D Gaussian Splatting [kerbl3Dgaussians] (3DGS) uses 3D Gaussians primitives to reconstruct a 3D scene from a set of images. Each Gaussian is defined with a position $\mu$, a rotation matrix $R$ (parameterized as a quaternion), scales $s\in\mathbb{R}^{3}$, opacity $o$ and view-dependent color $c$ as spherical harmonics coefficients. For rendering, Gaussians are projected to screen space and alpha-blended front-to-back, yielding the final color:
@@
2D Gaussian Splatting [2DGS] (2DGS) extends this by using flat 2D surfels better suited for reconstructing high-fidelity surfaces. Each gaussian defines a local tangent plane in world space, which is parameterized as:
@@ 3.3 Mesh Binding
Similarly to existing Gaussian head avatars [gaussianavatars, lee2025surfhead, HRAvatar, flashavatar], we bind Gaussians to triangles of the FLAME mesh. Within each triangle, $n$ splats are initialized with random barycentric coordinates within $[0,1]$. Each Gaussian has a learnable rotation $r$ relative to the orientation of its triangle (represented as a quaternion), a displacement $d\in\mathbb{R}$ along the triangle normal, scales $s\in\mathbb{R}^{2}$, and opacity $o\in\mathbb{R}$.
Given pose and expression parameters for FLAME, we first compute the deformed vertex positions $V_{d}\in\mathbb{R}^{3V}$ and face normals $N_{d}\in\mathbb{R}^{3F}$. The position of each Gaussian is obtained by interpolating the deformed vertex positions of its triangle with barycentric coordinates. The rotation is the Gaussian’s relative rotation $r$ multiplied by the triangle’s orientation matrix, whose columns are a tangent, bi-tangent and normal vector of the triangle plane.

### v2

@@ Abstract
Recent advancements in Gaussian Splatting have enabled increasingly accurate reconstruction of photorealistic head avatars, opening the door to numerous applications in visual effects, videoconferencing, and virtual reality. This, however, comes with the lack of intuitive editability offered by traditional triangle mesh-based methods. In contrast, we propose a method that combines the accuracy and fidelity of 2D Gaussian Splatting with the intuitiveness of UV texture mapping. By embedding each canonical Gaussian primitive’s local frame into a patch in the UV space of a template mesh in a computationally efficient manner, we reconstruct continuous editable material head textures from a single monocular video on a conventional UV domain. Furthermore, we leverage an efficient physically based reflectance model to enable relighting and editing of these intrinsic material maps. Through extensive comparisons with state-of-the-art methods, we demonstrate the accuracy of our reconstructions, the quality of our relighting results, and the ability to provide intuitive controls for modifying an avatar’s appearance and geometry via texture mapping without additional optimization.
[…]
volume: 45issue: 2
@@ 1 Introduction
Gaussian Splatting based avatars have revolutionized the capture and rendering of digital humans by delivering an unprecedented level of photorealism with real-time rendering capability, enabling new possibilities for reanimating content from video inputs. However, this realism comes with a trade-off: unlike traditional texture-map-based modeling, Gaussian avatars offer little flexibility for intuitive appearance editing. This gap becomes critical in practice. In film and visual effects, artists routinely adjust the smallest details of a face – smoothing skin to create a flawless appearance, removing distracting high-frequency features, or sculpting age and character through wrinkles, scars and bruises. In gaming and virtual production, creators seek the same level of control to personalize avatars with tattoos, makeup, or stylized patterns using artist-friendly tools that operate directly on material texture maps. Without such editing capabilities, even the most realistic avatar remains a fixed reproduction, rather than a medium for creative expression.
This limitation stems from the fact that each Gaussian splat carries its own color or material properties in isolation, without the shared structure that texture maps naturally provide. As a result, the surface lacks the local coherence needed to treat it as a continuous canvas, making it difficult – if not impossible – to apply edits consistently, whether to a small region of the face or to its overall appearance. Conversely, while mesh-based 3D inverse rendering methods [52, 4, 3] inherit a coherent defacto UV-domain fit for artist-friendly edits, they can suffer from topological rigidity when representing high-frequency details and are hampered in their reconstruction capacity due to the limitations of differentiable rasterization when handling complex or translucent geometries (cf. Table 1). Furthermore, naively embedding Gaussians in UV domain leads to discontinuous texture maps that are hard to edit successfully, as witnessed in our results (e.g. Figure 9) and also in [102]. The authors of FATE [102] proposed a second U-Net neural baking stage to alleviate this issue in the context of non-relightable avatars.
@@
Following seminal work [60, 42], we adopt FLAME-anchored [44] Gaussian splats. Our key observation is that, unlike vanilla 3DGS [35], which operates directly in screen space, the 2DGS variant [31] defines a local splat coordinate frame and computes ray-splat intersections to query kernels. We leverage this property to map ray-splat intersections into the UV domain via an approximate orthographic projection from the splat tangent space to its corresponding FLAME mesh triangle, effectively mapping each splat plane to a continuous UV patch. We devise an efficient method to compute this mapping using a single matrix multiplication per intersection, which is highly important to maintain the method’s efficiency given the very high number of intersections involved in rendering. PBR material attributes are then sampled smoothly from learnable texture maps (albedo, roughness, and specular reflectance) using bilinear filtering. The splat orientations are combined with a residual normal map to produce the final shading normals. We jointly optimize the FLAME parameters, texture maps, 2D Gaussian positional parameters and environment lighting. This enables differentiable expression of outgoing radiance via the Cook-Torrance BRDF [15] under the split-sum approximation [34], using splatted geometry and material G-buffers in a deferred shading framework. Finally, we introduce a novel UV regularization that is crucial for maintaining representation integrity by enforcing alignment of UV coordinates of ray-splat intersections for a given ray. Figure 9 underlines the benefits of our UV mapping and regularization compared to a naive UV embedding strategy based on the projection of splat origins.
@@
  • Expressive Gaussian primitives. Spatially varying attributes and normals, in a Phong shading [58] fashion, as opposed to single values, yield higher reconstruction quality with a smaller total model size than state-of-the-art relightable avatar competition [101] (Figure 6b). While similar expressiveness can also be achieved with small per-primitive textures (e.g. [64, 76, 9, 96, 88, 83, 71]), our approach uses a compact global map, offering both efficiency and editability.
@@
    • Intrinsic material supervision, such as our albedo regularization with the FLAME albedo model that guides physical decomposition mid-training and reduces artifacts (Figure 9). We note that we also incorporate a diffusion-based albedo prior in screen space, similar to [101], to further disambiguate the decomposition.
@@
We demonstrate the benefits of our method compared to traditional splat-based reconstruction and rendering techniques, and illustrate its versatility through various texture editing examples. Our method produces highly realistic head reconstructions that not only match or surpass the quality of state-of-the-art monocular relightable avatar HRAvatar [101] – despite its lack of explicit texture representations – but also unlocks new possibilities for artistic control.
@@ 2 Related work
Avatar Representations. The prior success of NeRFs [49, 51] brought impressive levels of photorealism in neural avatars based on differentiable volume rendering [30, 23, 29]. Rasterization approaches can alleviate the computational rendering burden of implicit neural ray casting. However, mesh-based rasterization [52] can yield subpar photometric reconstruction performance (e.g. FLARE [4] in Table 1) due to the limitations of meshes in modeling transparent and complex geometry. The recent advent of Gaussian Splatting [35] has enabled state-of-the-art reconstruction quality with real-time rendering [90, 98, 60]. Gaussian Splatting extends the EWA volume resampling framework [111, 110] to learnable [38] inverse rendering, modeling 3D scenes with explicit anisotropic Gaussian kernel primitives that can be sorted and rasterized efficiently in tiles. In this respect, the 2DGS variant [31] leverages planar 2D primitives instead of volumetric ones, and performs precise 2D kernel evaluation in object space as opposed to approximate ones in screen space (3DGS), thus leading to superior geometry and multi-view consistency. The superior normals provided by 2DGS [42, 67, 32, 16] have shown to be beneficial in physics-based inverse rendering [93, 96, 74, 77, 27] where inferring precise reflection directions [79] is pivotal. 2DGS also defines a local splat coordinate frame that allows us to map ray-splat intersections to our canonical UV space continuously, which is key in our novel avatar representation. Generalizable models for 2DGS [8, 33] have been recently proposed similarly to 3DGS [10, 13] before.
@@
Monocular & Relightable Avatars. Learning avatars from casual consumer grade videos paves the road to democratizing volumetric capture, in contrast to the costly requirements of light stages or multi-view setups [19, 28, 91]. Recent advances enable avatar learning from monocular video input (e.g. [106, 86, 14, 69, 46]), thanks to built-in inductive biases and robust monocular facial tracking (e.g. [20, 17, 63, 108]) and reconstruction (e.g. [26, 37, 25, 21, 22]) methods. Rigging avatars with underlying 3D morphable models (3DMMs) [44, 57, 2, 6, 61, 78, 54] enables efficient and robust learning and parametric animation control. They define a canonical space for implicit and explicit representations and factor out expression and pose deformations. Several works leverage data captured under controlled lighting conditions to produce relightable head avatars [65, 92] with dynamic radiance fields and physical reflectance models. Relightable Gaussian Codec Avatar [65] has shown good relighting capabilities for full multi-view OLAT captures through its material parameterization, and has also been extended to the phone scan setting [43]. Despite the increased ill-posedness of the problem, physics-based inverse rendering and relighting from mere monocular videos or sparse images is possible through further regularizations or appearance priors, such as modeling reflectance via simplified bidirectional reflectance distribution functions (BRDF) [4, 3], or relying on large data corpora [43]. Closest to our context, HRAvatar [101] employs deferred shading of FLAME-rigged 3D Gaussians augmented with material attributes. In contrast to the latter, we build on 2DGS deferred shading with the aim of extending controllability beyond relighting and reenactment, to additionally include intuitive local editing on conformal material maps, in the interest of more expressive and user-controllable avatar manipulation without sacrificing performance.
@@
UV Mapping & Editing. Recent research has sought to combine explicit surface parameterization with neural representations. Within the NeRF [49] framework, works such as NeuTex [85], Neural Gauge Field [100] and Nuvo [73] learn neural UV mappings that establish bijective correspondences between surfaces and textures, enabling surface-aware rendering across general or category-specific domains. Simultaneously, Gaussian Splatting can support 3D domain editing through semantic grouping and manipulation [94], dynamic 4D content [97], text-driven scene modifications [12], or brush painting [55]. Editing via screen-space gradients [70, 80, 24], which relies mostly on extensive optimization with pretrained diffusion models [103, 104] such as Instruct-Pix2Pix [7], suffers from high computational costs and lacks precise controllability. UV-domain editing requires continuous texture maps. While some methods structure Gaussians in the UV domain [39, 1, 69, 86, 80], most still operate with discontinuous sparse maps. Texture-GS [89] learns a neural UV mapping for 3DGS primitives for static objects, whilst we aim to use a predefined UV map for semantically grounded edits of a dynamic avatar. Contemporary work FATE [102] attempts to alleviate the texture map continuity issue for non-relightable dynamic Gaussian head avatars. However, it requires a separate learning stage involving neural baking via additional networks to obtain smooth editable textures and fails to preserve sharp details when edited, as shown in Figure 11. Our method avoids these multiple stages, bypassing the need for any neural UV mapping or neural baking. Moreover, to our knowledge, our approach is the first Gaussian head avatar to conjointly enable relighting and UV editing of appearance and normals.
2D Image-based relighting. A significant area of research addresses portrait relighting from monocular inputs directly, bypassing the need for full 3D reconstruction. Many approaches decompose the input image into various intrinsic components such as albedo, normals and specular reflectance [81, 66, 68, 40, 56], leveraging priors for accurate lighting decomposition. These priors often include parametric face models [5, 44, 2, 6, 61, 78, 54], light-stage datasets [18] or simplified physical models [58, 15]. In recent years, many works have used encoder-decoder architectures [48, 75, 107, 82, 53, 56, 62] or diffusion models [59, 99, 11] that can directly synthesize relit images. While these methods can capture complex illumination effects (e.g. hard shadows, inter-reflections and subsurface scattering), disentanglement of lighting from other factors remains difficult in the under-constrained monocular setting. Additionally, the lack of explicit 3D representations severely limits their ability to render novel viewpoints.
@@ 3.1 FLAME
The FLAME [44] 3D morphable model provides a template head mesh with joints for eyes, jaw and neck, and a statistical model for identity and expressions. Identity and expression deformations are expressed as blendshapes while the joints transform vertices through Linear Blend Skinning [47]. For simplicity, we refer to FLAME as template vertex positions $V_{t}\in\mathbb{R}^{3V}$, triangle topology $0..V-1]^{3F}$ and a deformation function $\mathcal{F}$:
@@
where the identity parameters are fixed and $\Psi$ is the concatenation of pose and expression parameters. We refer the reader to [44] for exhaustive details on this 3DMM. Additionally, FLAME provides per-corner UV coordinates for texturing (i.e. UV parameterization for the unwrapped template mesh).
@@ 3.2 Gaussian Splatting
3D Gaussian Splatting [36] (3DGS) uses 3D Gaussians primitives to reconstruct a 3D scene from a set of images. Each Gaussian is defined with a position $\mu$, a rotation matrix $R$ (parameterized as a quaternion), scales $s\in\mathbb{R}^{3}$, opacity $o$ and view-dependent color $c$ as spherical harmonics coefficients. For rendering, Gaussians are projected to screen space and alpha-blended front-to-back, yielding the final color:
@@
2D Gaussian Splatting [31] (2DGS) extends this by using flat 2D surfels better suited for reconstructing high-fidelity surfaces. Each gaussian defines a local tangent plane in world space, which is parameterized as:
@@ 3.3 Mesh Binding
Representation. Similarly to existing Gaussian head avatars [60, 41, 101, 86], we bind Gaussians to triangles of the FLAME mesh. Within each triangle, $n$ splats are initialized with random barycentric coordinates within $[0,1]$. Each Gaussian has a learnable rotation $r$ relative to the orientation of its triangle (represented as a quaternion), a displacement $d\in\mathbb{R}$ along the triangle normal, scales $s\in\mathbb{R}^{2}$, and opacity $o\in[0,1]$.
Deformation. Given pose and expression parameters for FLAME, we first compute the deformed vertex positions $V_{d}\in\mathbb{R}^{3V}$ and face normals $N_{d}\in\mathbb{R}^{3F}$. The position of each Gaussian is obtained by interpolating the deformed vertex positions of its triangle with barycentric coordinates. The rotation is the Gaussian’s relative rotation $r$ multiplied by the triangle’s orientation matrix, whose columns are a tangent, bi-tangent and normal vector of the triangle plane.